# Neural Networks and Friction: Slide, Hold, Learn

### 23 Feb 2024 v2

@@ I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3, 4]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [5].
@@
Two notable advancements in RNN architecture are the Gated Recurrent Unit (GRU) and Long Short-Term Memory (LSTM) networks. Both GRU and LSTM address a fundamental limitation of basic RNNs known as the vanishing gradient problem, which makes it difficult for RNNs to learn dependencies between events that occur at long intervals in the input data [30]. GRUs simplify the LSTM architecture while retaining its ability to capture long-term dependencies, making it both efficient and powerful for a wide range of tasks [31].
The application of RNNs, GRUs, and LSTMs has been widespread and transformative across many fields, including physics [32]. Particularly, their success in modeling path-dependent plasticity of materials highlights their potential [33, 34, 35, 36, 37]. Path-dependent plasticity is a critical aspect in the study of materials science, where the history of material deformation affects its current and future state. RNNs have been shown to effectively model these complex relationships, capturing the history-dependent nature of material behavior [38, 39].
@@
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws, the aging law in particular, eq. 2. This approach is essentially different from the one used recently by Ishiyama and collaborators [40], who used machine learning (random forest [41]) to improve the estimation of RSF parameters eq. 3: machine learning is not deployed to improve an empirical approximate law, but to replace it altogether, down the road, in larger-scale simulations[16, 42, 43]. It also differs from the use made by Rouet-Leduc et al [44] and Hulbert et al [45], among others [46, 47], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. Herein, we assume a virtual interface that is perfectly modeled by a traditional slip-law eqs. 3 and 2, for which data for training and verification can be generated easily. We use those data to train a custom GRU network that learns the dynamics without internal variables, since the state is not a physical quantity that can be measured experimentally (assuming it even exists as such). In reality, the complexities of interface physics [48] are such that probably no simplified internal variable can be defined as long as one aims at capturing any modestly ambitious degree of complexity. We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the parameters of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated by the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied.
@@ II.3.1 Architecture
The main net, which is aimed at replacing eq. 3 and whose parameters must be learned via training, is made up by a GatedRecurrentLayer, followed by a folded LinearLayer to transforms the output of the GRU through a linear transformation.
@@ II.3.3 Loss function
Finally, interface strengthening during hold periods (a.k.a. healing) is recognized to happen logarithmically with respect to hold time, i.e., counting from the moment that the interface ceases to move. Consider that, given a sequence, the hold period starts at the $j$-th entry, and it lasts until $(j+N_{h})$-th entry. We can gather these terms in a separate vector $\delta\bm{\mu}_{h}=\left\{\delta\mu_{i}\right\}_{i=k}^{i=k+N_{h}}=\left\{\delta\mu_{n}\right\}_{n=1}^{n=N_{h}}$, where the sub-index $n$ has been introduced to index the values during hold straightforwardly. So the healing condition can be expressed as:
@@ VI Final Remarks
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several promising avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the exploration of alternative RNN architectures and more nuanced physics-informed loss functions (for instance, to account for friction manifesting indirectly in the dynamics of spring-block systems [2]). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of volumes of available data: unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [54, 55], aimed at improving data economy in the context of constitutive modeling with neural networks, can help in this case.
As mentioned above, incorporating spring-block dynamics into the model could pave the way for learning from real experimental data, thus circumventing the gap between theoretical models and practical observations. Combined with the use of automatic differentiation [53] (not explored at this point), it could offer a completely new pathway to take advantage of laboratory restuls to understand friction, replacing the use of experiments to simply fit parameters by producing training data that is the infused in a network that can be used later for predictions.
@@ Acknowledgements
The author acknowledges financial support from the Swiss National Science Foundation, via Ambizione Grant PZ00P2_216341 “Data-Driven Computational Friction”. Useful conversations (regarding text improvements, rate-and-state friction and/or laboratory experiments) with Dr. Gabriel Meyer, Dr. Simon Guérin-Marthe and Ms. Roxane Ferry are gratefully acknowledged. An EPFL bachelor project conducted by Mr. Antoine Binggeli provided Tensorflow code that inspired the use of GRUs to this particular problem.

### 28 Feb 2024 v3

@@ I Introduction
Understanding the dynamics of friction is paramount across several disciplines, including mechanical engineering, civil engineering, and geophysics [1, 2]. In mechanical and civil engineering, precise knowledge of frictional forces is crucial for the design and analysis of machines, structures, and materials under varying operational conditions [3, 4, 5]. In geophysics, friction dynamics play a critical role in earthquake mechanics, influencing the initiation, propagation, and arrest of seismic slips along faults [6].
@@
Two notable advancements in RNN architecture are the Gated Recurrent Unit (GRU) and Long Short-Term Memory (LSTM) networks. Both GRU and LSTM address a fundamental limitation of basic RNNs known as the vanishing gradient problem, which makes it difficult for RNNs to learn dependencies between events that occur at long intervals in the input data [33, 31]. GRUs simplify the LSTM architecture while retaining its ability to capture long-term dependencies, making it both efficient and powerful for a wide range of tasks [32].
The application of RNNs, GRUs, and LSTMs has been widespread and transformative across many fields, including physics [34]. Particularly, their success in modeling path-dependent plasticity of materials highlights their potential [35, 36, 37, 38, 39, 40, 41, 42]. Path-dependent plasticity is a critical aspect in the study of materials science, where the history of material deformation affects its current and future state. RNNs have been shown to effectively model these complex relationships, capturing the history-dependent nature of material behavior [43, 44].
@@
The goal of this short communication is to show that RNNs, based on GRUs, can perform at least as well as RSF laws, the aging law in particular, eq. 2. This approach is essentially different from the one used recently by Ishiyama and collaborators [45], who used machine learning (random forest [46]) to improve the estimation of RSF parameters eq. 3: machine learning is not deployed to improve an empirical approximate law, but to replace it altogether, down the road, in larger-scale simulations[17, 47, 48]. It also differs from the use made by Rouet-Leduc et al [49] and Hulbert et al [50], among others [51, 52], who used machine learning techniques to devise impending rupture predictions based on acoustic emissions recorded in lab settings. To the best of our knowledge, the closest work to ours may be the one by Wang and Sun[53], which focused on interfacial traction-opening laws, and the recent one by Cravero and Bhattacharya [54], who studied interface adhesion. Herein, we assume a virtual interface that is perfectly modeled by a traditional slip-law eqs. 3 and 2, for which data for training and verification can be generated easily. We use those data to train a custom GRU network that learns the dynamics without explicitly defining internal variables [43], since the state is not a physical quantity that can be measured experimentally, cf. work by Kumar and collaborators [55, 56]. In reality, the complexities of interface physics [57] are such that probably no simplified internal variable can be defined as long as one aims at capturing any modestly ambitious degree of complexity. We presume implicitly that sliding amount and sliding rate are the sole driving forces of the phenomenon, and whatever internal variables there may be can be internally codified in the parameters of the net. The figure of merit will not be the dynamic friction coefficient itself, but variations in the friction coefficient due to sliding. This decision is motivated by the practical inability to surmise the time evolution of the friction coefficient except if proper initial conditions are supplied.
@@ II.3.1 Architecture
The main net, which is aimed at replacing eq. 3 and whose parameters must be learned via training, is made up by a single GatedRecurrentLayer, followed by a folded LinearLayer to transforms the output of the GRU through a linear transformation.
@@ II.3.3 Loss function
Finally, interface strengthening during hold periods (a.k.a. healing) is recognized to happen logarithmically with respect to hold time, i.e., counting from the moment that the interface ceases to move. Consider that, given a sequence, the hold period starts at the $j$-th entry, and it lasts until $(j+N_{h})$-th entry. We can gather these terms in a separate vector $\delta\bm{\mu}_{h}=\left\{\delta\mu_{i}\right\}_{i=k}^{i=k+N_{h}}=\left\{\delta\mu_{n}\right\}_{n=1}^{n=N_{h}}$, where the sub-index $n$ has been introduced to index the values during hold straightforwardly. Recall that the sampling time step is constant, then we can replace working with time by working with indices. So the healing condition can be expressed as:
@@ VI Final Remarks
This document contains but an exploratory approach to modeling frictional interface via deep learning. Looking forward, there are several promising avenues for enhancing the predictive power and generalizability of recurrent neural networks for dynamic friction. Improvement could be sought through the addition of more nuanced physics-informed loss functions predicated on thermodynamic principles [39, 42] and exploration of alternative or ad hoc[35, 43] RNN architectures). Further parametric studies in terms of network hyperparameters are also warranted, in particular in terms of network depth (stacking GRUs in different ways[35]) and the volumes of available data: unlike cheap synthetic data, experimental data remains relatively scarce; hence it is worth exploring how new techniques [63, 64], aimed at improving data economy in the context of constitutive modeling with neural networks, can help in this case.
Incorporating spring-block dynamics [2] into the model could pave the way for learning from real experimental data, thus circumventing the gap between theoretical models and practical observations. Combined with the use of automatic differentiation [62] (not explored at this point), it could offer a completely new pathway to take advantage of laboratory results to understand friction, replacing the use of experiments to simply fit parameters by producing training data that is the infused in a network that can be used later for predictions.
@@ Acknowledgements
The author acknowledges financial support from the Swiss National Science Foundation, via Ambizione Grant PZ00P2_216341 “Data-Driven Computational Friction”. Useful conversations (regarding text improvements, rate-and-state friction and/or laboratory experiments) with Dr. Gabriel Meyer, Dr. Simon Guérin-Marthe and Ms. Roxane Ferry are gratefully acknowledged. An EPFL bachelor project conducted by Mr. Antoine Binggeli developed Tensorflow code that inspired the use of GRUs to this particular problem.